In the 1992 presidential election, Alaska's 40 election districts averaged 2118 votes per district for President Clinton. The standard deviation was 577. (There are only 40 election districts in Alaska.) The distribution of the votes per district for President Clinton was bell-shaped. Let X = number of votes for President Clinton for an election district. (Source: The World Almanac and Book of Facts) Round all answers except part e. to 4 decimal places. a. What is the distribution of X? X ~ N(?,?) b. Is 2118 a population mean or a sample mean? c. Find the probability that a randomly selected district had fewer than 2005 votes for President Clinton. d. Find the probability that a randomly selected district had between 2236 and 2434 votes for President Clinton. e. Find the first quartile for votes for President Clinton. Round your answer to the nearest whole number.
Answer #1

Let x = The number of votes for President Clinton for on election distoict. a) The distribution of x.is XnNcmean = 2118, J =d) Pf 2236 5x52434) = P(x)2236) - pfxd 2434) = D( 854 ) 2236-2118 599 -Pfx-4) 2434-2118 r 599 = 04219 0. 2989 0:12.30 The pro